OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there?

OpenStudy (johnweldon1993):

Well...you know that both only have 1 head...so let human = h and horse = r h + r = 74 humans have 2 legs.....horses have 4 legs...so 2h + 4r = 196 solve the system

OpenStudy (johnweldon1993):

Just to write em both together h + r = 74 2h + 4r = 196 Solve the system of equations...hope that makes sense